OpenStudy (anonymous):

Write the slope-intercept form of the equation of the line parallel to y = 5x + 2, which passes through the point (4, 9).

OpenStudy (hoblos):

parallel lines have same slope so slope = 5 y-9=5(x-4) y= 5x-11